# Hulls of projective Reed-Muller codes over the projective plane

### v2

@@ 1. Introduction
Going back to projective Reed-Muller codes, as they are evaluation codes, we can view their codewords as classes of polynomials in a quotient ring. This motivates Section 3 of this paper, where we compute bases of polynomials for some appropriate subspaces of the quotient rings associated to projective Reed-Muller codes by using Gröbner bases techniques. As a consequence of this computation, we give a basis for the relative hull of projective Reed-Muller codes over the projective plane $\mathbb{P}^{2}$. We also estimate the dimension of the Hermitian hull, and give bases in some cases. The estimate that we obtain is sharp in all the cases we have checked. As a byproduct of these computations, we also compute the Hermitian hull of affine Reed-Muller codes in 2 variables, which was known to be trivial in some cases [11, 22], but was not known in general. With this knowledge, in Section 4 we give the parameters of the EAQECCs obtained by using projective Reed-Muller codes over $\mathbb{P}^{2}$, since the dimension of the hull determines the parameter $c$. Although projective Reed-Muller codes had already been used to construct QECCs in some particular cases [23], the cases that required entanglement assistance had not been yet addressed. We focus on the case of $\mathbb{P}^{2}$ because it offers a good trade-off between providing long codes over a small finite field and avoiding computations that are too involved, making explicit formulas unfeasible.
@@ 2. Preliminaries
The Reed-Muller code $\operatorname{RM}_{d}(q,m)$, $1\leq d\leq m(q-1)$, is an $[n,k]$-code with
@@
In what follows, we will abuse the notation and denote $S_{d}/I(P^{m})=(S_{d}+I(P^{m}))/I(P^{m})$. Moreover, for a polynomial $f\in S$, we will use the same notation $f$ for both the polynomial and its class in $S/I(P^{m})$. We have the following result for the vanishing ideal of $P^{m}$ from [13].
@@
Moreover, these generators form a universal Gröbner basis of the ideal $I(P^{m})$, and we have that
@@
Remark 2.7.
Any monomial $x_{0}^{\alpha_{0}}x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}$ with $\alpha_{0}>0$ is equivalent to $x_{0}^{\alpha^{\prime}_{0}}x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}$ in $S/I(P^{2})$, for any $\alpha^{\prime}_{0}>0$, because we have the polynomial $x_{0}^{2}-x_{0}$ in $I(P^{2})$. In particular, $x_{0}^{\alpha_{0}}x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}$ is equivalent to $x_{0}x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}$ if $\alpha_{0}>0$.
@@
From [13], we have the following lemmas about $S/I(P^{2})$.
@@
Let $a_{0},a_{1},a_{2}$ be integers, with $a_{0}>0$, $a_{1}>0$. We have that
$x_{0}^{a_{0}}x_{1}^{a_{1}}x_{2}^{a_{2}}$ $\equiv x_{1}^{a_{1}}x_{2}^{a_{2}}+x_{0}x_{2}^{a_{2}}-x_{2}^{a_{2}}+x_{0}x_{1}-x_{0}-x_{1}+1\bmod I(P^{2})$
$\equiv x_{1}^{a_{1}}x_{2}^{a_{2}}+(x_{0}-1)(x_{2}^{a_{2}}+x_{1}-1)\bmod I(P^{2}).$
Lemma 2.11 allows us to express all the monomials from the basis in Lemma 2.8 in terms of the basis from Lemma 2.10. This is crucial because the idea of the next section is, for $1\leq d_{1}\leq d_{2}\leq 2(q-1)$, to consider the bases from Lemma 2.8 for $S_{d_{1}}/I(P^{2})$ and $S_{d_{2}}/I(P^{2})$, and express them in terms of the basis for $S/I(P^{2})$ from Lemma 2.10 using Lemma 2.11. Once we have all the polynomials expressed in this way, it is easier to find the polynomials lying in $S_{d_{1}}/I(P^{2})\cap S_{d_{2}}/I(P^{2})$, which, as we will see, determines the relative hull of $\operatorname{PRM}_{d_{1}}(2)$ and $\operatorname{PRM}_{d_{2}}(2)$. For the Hermitian case, the ideas are also very similar, although the computations are more involved.
@@ 3.1. Euclidean hull
In this subsection we compute a basis for $\text{Hull}_{C_{2}}(C_{1})=C_{1}\cap C_{2}^{\perp}$, the relative hull of $C_{1}$ and $C_{2}$, when $C_{i}=\operatorname{PRM}_{d_{i}}(2)$, for $i=1,2$. For $\operatorname{PRM}_{d_{2}}^{\perp}(2)$, by Theorem 2.2 we have that, if $d_{2}\not=q-1$, then $\operatorname{PRM}_{d_{2}}^{\perp}(2)=\operatorname{PRM}_{d_{2}^{\perp}}(2)$, where $d_{2}^{\perp}=2(q-1)-d_{2}$ (assuming $\operatorname{PRM}_{0}(2)=\langle(1,\dots,1)\rangle$). We avoid the case where $d_{2}=q-1$ because, by Theorem 2.2, in that case we have $\operatorname{PRM}_{d_{2}}^{\perp}(2)=\operatorname{PRM}_{d_{2}^{\perp}}(2)+\langle(1,\dots,1)\rangle$, which is no longer isomorphic to $S_{d_{2}^{\perp}}/I(P^{2})$. Assuming $d_{2}\not=q-1$, the problem of obtaining a basis for $\operatorname{PRM}_{d_{1}}(2)\cap\operatorname{PRM}_{d_{2}}^{\perp}(2)$ becomes equivalent to computing a basis for $\operatorname{PRM}_{d_{1}}(2)\cap\operatorname{PRM}_{d_{2}}(2)$, for any $1\leq d_{1}\leq d_{2}\leq 2(q-1)$, $d_{2}\not=q-1$, and that is the problem we solve in what follows. In [23, Thm. 10.7] the authors construct quantum codes using pairs of projective Reed-Muller codes such that the dual of one of the codes is contained in the other. In particular, they obtain the following result.
@@
In the following, for an integer $0\leq z\leq 2(q-1)$, we denote by $\overline{z}$ the integer $11\leq\overline{z}\leq q-$ 1 such that $\overline{z}\equiv z\bmod q-1$ if $z>0$, and $\overline{z}=0$ if $z=0$. This ensures that we always have
for any $0\leq a_{0},a_{1},a_{2}\leq 2(q-1)$.
@@
We have seen that $x_{0}^{q-1}x_{1}^{d_{2}-d_{1}-(q-1)}x_{2}^{d_{1}}$ is the only monomial that we can consider to obtain the monomial $x_{2}^{d_{1}}$ in the right hand side of (5), and we need to consider $x_{1}^{d_{2}-d_{1}}x_{2}^{d_{1}}$ with the opposite coefficient to cancel the monomial $x_{1}^{d_{2}-d_{1}-(q-1)}x_{2}^{d_{1}}$ (the expression of $x_{1}^{d_{2}-d_{1}}x_{2}^{d_{1}}$ in terms of the basis from Lemma 2.10), i.e., we must have $-\gamma_{q-1,d_{2}-d_{1}-(q-1),d_{1}}=\gamma_{0,d_{2}-d_{1},d_{1}}=\lambda_{0,d_{1}}$. We can define in this case $(A_{2}^{d_{2}})^{(1)}=A_{2}^{d_{2}}\setminus\{x_{0}^{q-1}x_{1}^{d_{2}-d_{1}-(q-1)}x_{2}^{d_{1}},x_{1}^{d_{2}-d_{1}}x_{2}^{d_{1}}\}$, and consider
@@ 3.2. Hermitian hull
The Hermitian dual of a code $C\subset{\mathbb{F}}_{q^{2}}^{n}$ is defined as $C^{\perp_{h}}:=\{v\in{\mathbb{F}}_{q^{2}}^{n}\mid v\cdot_{h}w=0,\;\forall\;w\in C\}$. We recall that we defined the Hermitian hull as $\text{Hull}^{H}(C)=C\cap C^{\perp_{h}}$. It is easy to check that, for a code $C\subset{\mathbb{F}}_{q^{2}}^{n}$, we have that $C^{\perp_{h}}=(C^{\perp})^{q}$, where we consider the component wise power of $q$. In particular, this implies that the Hermitian dual and the Euclidean dual have the same parameters. In this section we show that we may apply similar techniques to the ones used in the previous section to compute the Hermitian hull in some cases. In what follows, as we are working over ${\mathbb{F}}_{q^{2}}$, we change $q$ by $q^{2}$ in the definitions of $S$, $A_{i}^{d}$, for $i=1,2,3$, etc. For instance, $S={\mathbb{F}}_{q^{2}}[x_{0},x_{1},x_{2}]$ in this section.
@@
The codes’ inclusion $\operatorname{RM}_{d}(q^{2},m)\subset\operatorname{RM}_{d}^{\perp_{h}}(q^{2},m)$ holds if, and only if, $11\leq d\leq m(q-1)-$ 1.
@@
The image by the evaluation map over $\mathbb{A}^{2}$ of the set of monomials $U_{d_{1},d_{2}}=\{x_{1}^{a_{1}}x_{2}^{a_{2}}\mid 0\leq a_{1},a_{2}\leq q^{2}-1,\;a_{1}+a_{2}\leq d_{1},\;\overline{qa_{1}}+\overline{qa_{2}}\leq 2(q^{2}-1)-d_{2}-1\}$ is a basis for $\operatorname{RM}_{d_{1}}(q^{2},m)\cap\operatorname{RM}_{d_{2}}^{\perp_{h}}(q^{2},m)$.
@@
The monomials $x_{1}^{a_{1}}x_{2}^{a_{2}}$ with $10\leq a_{1},a_{2}\leq q^{2}-$ 1 have linearly independent evaluations over $\mathbb{A}^{2}$, and their evaluations generate ${\mathbb{F}}_{q^{2}}^{2}$. The evaluation of a monomial $x_{1}^{a_{1}}x_{2}^{a_{2}}$ with $10\leq a_{1},a_{2}\leq q^{2}-$ 1 and $a_{1}+a_{2}\leq d_{1}$ is in $\operatorname{RM}_{d_{2}}^{\perp_{h}}(q^{2},2)$ if and only if $x_{1}^{a_{1}}x_{2}^{a_{2}}\equiv(x_{1}^{b_{1}}x_{2}^{b_{2}})^{q}\bmod I(\mathbb{A}^{2})$ for some $10\leq b_{1},b_{2}\leq q^{2}-$ 1 such that $b_{1}+b_{2}\leq 2(q^{2}-1)-d_{2}-1$, where $I(\mathbb{A}^{2})=\langle x_{1}^{q^{2}}-x_{1},x_{2}^{q^{2}}-x_{2}\rangle$ (we have used the duality from Theorem 2.5 and the fact that $\operatorname{RM}_{d_{2}}^{\perp_{h}}(q^{2},2)=(\operatorname{RM}_{d_{2}}^{\perp}(q^{2},2))^{q}$). The previous happens if and only if $a_{i}\equiv qb_{i}\bmod q^{2}-1$, for $i=1,2$, which is equivalent in this case to having $b_{i}=\overline{qa_{i}}$, for $i=1,2$. ∎
@@
In this section we are only considering projective Reed-Muller codes over ${\mathbb{F}}_{q^{2}}$, and we understand by $\operatorname{PRM}_{d}(2)$ that we are considering the code $\operatorname{PRM}_{d}(q^{2},2)$. Using that $C^{\perp_{h}}=(C^{\perp})^{q}$, if we consider $(A_{i}^{d})^{q}:=\{(x^{\alpha})^{q}\mid x^{\alpha}\in A_{i}^{d}\}$ and $d^{\perp}=2(q^{2}-1)-d$, we have that the image by the evaluation map of $(A^{d^{\perp}})^{q}:=\bigcup_{i=1}^{3}(A_{i}^{d^{\perp}})^{q}$ is a basis for $\operatorname{PRM}_{d}^{\perp_{h}}(2)=(\operatorname{PRM}_{d^{\perp}}(2))^{q}$ (if $d\not\equiv 0\bmod q^{2}-1$). Following the notation from the previous section, we will denote by $S_{d^{\perp}}^{q}/I(P^{2})$ the vector space generated in $S/I(P^{2})$ by $\bigcup_{i=1}^{3}(A_{i}^{d^{\perp}})^{q}$.
@@
$\operatorname{PRM}_{d}(2)\cap\operatorname{PRM}_{d^{\perp}}^{q}(2)=(\operatorname{PRM}_{d}^{q}(2)\cap\operatorname{PRM}_{d^{\perp}}(2))^{q},$
and then at the right-hand side of the previous equality we have the Hermitian hull of a projective Reed-Muller code of degree $d^{\perp}<q^{2}-1$, to the power of $q$. Moreover, because of Theorem 2.2 we are going to avoid the case with $d=q^{2}-1$ when giving results for the Hermitian hull (for results about $S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$ we will still consider $d=q^{2}-1$). This is because in that case $\operatorname{PRM}_{d}^{\perp}(2)=\operatorname{PRM}_{d^{\perp}}(2)+\langle(1,\dots,1)\rangle\neq\operatorname{PRM}_{d^{\perp}}(2)$, and $\operatorname{PRM}_{d}^{\perp_{h}}(2)=(\operatorname{PRM}_{d^{\perp}}(2)+\langle(1,\dots,1)\rangle)^{q}\neq\operatorname{PRM}_{d^{\perp}}^{q}(2)\cong S_{d^{\perp}}^{q}/I(P^{2})$. Hence, if $d=q^{2}-1$ we do not have the isomorphism between $\operatorname{PRM}_{d}^{\perp_{h}}(2)$ and $S_{d^{\perp}}^{q}/I(P^{2})$, and also this case is the least interesting for quantum codes because we do not have a bound for the minimum distance of the dual code.
@@
In the following results we argue in a similar way to Section 3.1 to show which polynomials can be in $S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$ depending on whether the monomials in the support of these polynomials are contained in $A_{2}^{d}$ or if they have $x_{2}^{d}$ in their support (we recall that $A_{3}^{d}=\{x_{2}^{d}\}$). We restrict to the case $1\leq d\leq 2(q-1)$ in some results because in that case we have $U=A_{1}^{d}$ by Lemma 3.16, which is similar to what happens in the Euclidean case.

### v3

@@ 1. Introduction
Going back to projective Reed-Muller codes, as they are evaluation codes, we can view their codewords as classes of polynomials in a quotient ring. This motivates Section 3 of this paper, where we compute bases of polynomials for some appropriate subspaces of the quotient rings associated to projective Reed-Muller codes by using Gröbner bases techniques. As a consequence of this computation, we give a basis for the relative hull of projective Reed-Muller codes over the projective plane $\mathbb{P}^{2}$. We also estimate the dimension of the Hermitian hull, and give bases in some cases. The estimate that we obtain is sharp in all the cases we have checked. As a byproduct of these computations, we also compute the Hermitian hull of affine Reed-Muller codes in 2 variables, which was known to be trivial in some cases [12, 24], but was not known in general. With this knowledge, in Section 4 we give the parameters of the EAQECCs obtained by using projective Reed-Muller codes over $\mathbb{P}^{2}$, since the dimension of the hull determines the parameter $c$. Although projective Reed-Muller codes had already been used to construct QECCs in some particular cases [25], the cases that required entanglement assistance had not been yet addressed. We focus on the case of $\mathbb{P}^{2}$ because it offers a good trade-off between providing long codes over a small finite field and avoiding computations that are too involved, making explicit formulas unfeasible. For the general case of $\mathbb{P}^{m}$, one has to make additional assumptions, such as restricting to the Euclidean case and requiring $C_{1}=C_{2}$, see [17].
@@ 2. Preliminaries
The Reed-Muller code $\operatorname{RM}_{d}(q,m)$, $0\leq d\leq m(q-1)$, is an $[n,k]$-code with
@@
In what follows, we will abuse the notation and denote $S_{d}/I(P^{m})=(S_{d}+I(P^{m}))/I(P^{m})$. Moreover, for a polynomial $f\in S$, we will use the same notation $f$ for both the polynomial and its class in $S/I(P^{m})$. We refer the reader to [7] for the basic concepts of Gröbner bases. For $f\in S$, we denote by $\operatorname{in}(f)$ the leading monomial of $f$ (without the coefficient). For an ideal $I\subset S$, $\operatorname{in}(I)$ denotes the ideal generated by the leading monomials of the polynomials in $I$. We have the following result for the vanishing ideal of $P^{m}$ from [14].
@@
Moreover, these generators form a universal Gröbner basis of the ideal $I(P^{m})$ (i.e., they form a Gröbner basis for any monomial order), and we have that
@@
From [14], we have the following lemma about $S/I(P^{2})$.
@@
We show now how to express any monomial from $A^{d}$ in terms of the basis from Lemma 2.9. The following definition is useful for this purpose.
Definition 2.10.
For an integer $z\geq 0$, we denote by $\overline{z}$ the integer $11\leq\overline{z}\leq q-$ 1 such that $\overline{z}\equiv z\bmod q-1$ if $z>0$, and $\overline{z}=0$ if $z=0$.
Remark 2.11.
Any monomial $x_{0}^{\alpha_{0}}x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}$ with $\alpha_{0}>0$ is equivalent to $x_{0}^{\alpha^{\prime}_{0}}x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}$ in $S/I(P^{2})$, for any $\alpha^{\prime}_{0}>0$, because we have the polynomial $x_{0}^{2}-x_{0}$ in $I(P^{2})$. In particular, $x_{0}^{\alpha_{0}}x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}$ is equivalent to $x_{0}x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}$ if $\alpha_{0}>0$. Moreover, in Definition 2.10 we treat the case $z=0$ separately so that we have
for any $0\leq a_{0},a_{1},a_{2}\leq 2(q-1)$. Notice that, for evaluation codes, having exponent $q-1$ is not the same as 00 since, for instance,
$x_{0}^{0}=1\not\equiv x_{0}^{q-1}\bmod I(P^{2}).$
This can be checked by evaluating $1$ and $x_{0}^{q-1}$ at any point of the form $(0,1,x_{2})$, $x_{2}\in\mathbb{F}_{q}$ (recall that two polynomials are equivalent modulo $I(P^{2})$ if and only if they have the same evaluation).
The following Lemma from [14] shows how to express any monomial in terms of the basis from Lemma 2.9.
@@
Let $a_{0},a_{1},a_{2}$ be integers.
  1. (1)
    If $a_{0}=0$, then
    $x_{1}^{a_{1}}x_{2}^{a_{2}}\equiv x_{1}^{\overline{a_{1}}}x_{2}^{\overline{a_{2}}}\bmod I(P^{2}).$
  2. (2)
    If $a_{1}=0$, then
    $x_{0}^{a_{0}}x_{2}^{a_{2}}\equiv x_{0}x_{2}^{\overline{a_{2}}}\bmod I(P^{2}).$
  3. (3)
    If $a_{0}>0$ and $a_{1}>0$, then
    $x_{0}^{a_{0}}x_{1}^{a_{1}}x_{2}^{a_{2}}$ $\equiv x_{1}^{\overline{a_{1}}}x_{2}^{\overline{a_{2}}}+x_{0}x_{2}^{\overline{a_{2}}}-x_{2}^{\overline{a_{2}}}+x_{0}x_{1}-x_{0}-x_{1}+1\bmod I(P^{2})$
    $\equiv x_{1}^{\overline{a_{1}}}x_{2}^{\overline{a_{2}}}+(x_{0}-1)(x_{2}^{\overline{a_{2}}}+x_{1}-1)\bmod I(P^{2}).$
In particular, Lemma 2.12 allows us to express all the monomials from the basis in Lemma 2.7 in terms of the basis from Lemma 2.9. This is crucial because the idea of the next section is, for $1\leq d_{1}\leq d_{2}\leq 2(q-1)$, to consider the bases from Lemma 2.7 for $S_{d_{1}}/I(P^{2})$ and $S_{d_{2}}/I(P^{2})$, and express them in terms of the basis for $S/I(P^{2})$ from Lemma 2.9 using Lemma 2.12. Once we have all the polynomials expressed in this way, it is easier to find the polynomials lying in $S_{d_{1}}/I(P^{2})\cap S_{d_{2}}/I(P^{2})$, which, as we will see, determines the relative hull of $\operatorname{PRM}_{d_{1}}(2)$ and $\operatorname{PRM}_{d_{2}}(2)$. For the Hermitian case, the ideas are also very similar, although the computations are more involved.
@@ 3.1. Euclidean hull
In this subsection we compute a basis for $\text{Hull}_{C_{2}}(C_{1})=C_{1}\cap C_{2}^{\perp}$, the relative hull of $C_{1}$ and $C_{2}$, when $C_{i}=\operatorname{PRM}_{d_{i}}(2)$, for $i=1,2$. For $\operatorname{PRM}_{d_{2}}^{\perp}(2)$, by Theorem 2.2 we have that, if $d_{2}\not=q-1$, then $\operatorname{PRM}_{d_{2}}^{\perp}(2)=\operatorname{PRM}_{d_{2}^{\perp}}(2)$, where $d_{2}^{\perp}=2(q-1)-d_{2}$ (assuming $\operatorname{PRM}_{0}(2)=\langle(1,\dots,1)\rangle$). We avoid the case where $d_{2}=q-1$ because, by Theorem 2.2, in that case we have $\operatorname{PRM}_{d_{2}}^{\perp}(2)=\operatorname{PRM}_{d_{2}^{\perp}}(2)+\langle(1,\dots,1)\rangle$, which is no longer isomorphic to $S_{d_{2}^{\perp}}/I(P^{2})$. Assuming $d_{2}\not=q-1$, the problem of obtaining a basis for $\operatorname{PRM}_{d_{1}}(2)\cap\operatorname{PRM}_{d_{2}}^{\perp}(2)$ becomes equivalent to computing a basis for $\operatorname{PRM}_{d_{1}}(2)\cap\operatorname{PRM}_{d_{2}}(2)$, for any $1\leq d_{1}\leq d_{2}\leq 2(q-1)$, $d_{2}\not=q-1$, and that is the problem we solve in what follows. In [25, Thm. 10.7] the authors construct quantum codes using pairs of projective Reed-Muller codes such that the dual of one of the codes is contained in the other. In particular, they obtain the following result.
@@
We have seen that $x_{0}^{q-1}x_{1}^{d_{2}-d_{1}-(q-1)}x_{2}^{d_{1}}$ is the only monomial that we can consider to obtain the monomial $x_{2}^{d_{1}}$ in the right hand side of (5), and we need to consider $x_{1}^{d_{2}-d_{1}}x_{2}^{d_{1}}$ with the opposite coefficient to cancel the monomial $x_{1}^{d_{2}-d_{1}-(q-1)}x_{2}^{d_{1}}$ (the expression of $x_{1}^{d_{2}-d_{1}}x_{2}^{d_{1}}$ in terms of the basis from Lemma 2.9), i.e., we must have $-\gamma_{q-1,d_{2}-d_{1}-(q-1),d_{1}}=\gamma_{0,d_{2}-d_{1},d_{1}}=\lambda_{0,d_{1}}$. We can define in this case $(A^{d_{2}})^{(1)}=A^{d_{2}}\setminus\{x_{0}^{q-1}x_{1}^{d_{2}-d_{1}-(q-1)}x_{2}^{d_{1}},x_{1}^{d_{2}-d_{1}}x_{2}^{d_{1}}\}$, and consider
@@ 3.2. Hermitian hull
The Hermitian dual of a code $C\subset{\mathbb{F}}_{q^{2}}^{n}$ is defined as $C^{\perp_{h}}:=\{v\in{\mathbb{F}}_{q^{2}}^{n}\mid v\cdot_{h}w=0,\;\forall\;w\in C\}$. We recall that we defined the Hermitian hull as $\text{Hull}^{H}(C)=C\cap C^{\perp_{h}}$. It is easy to check that, for a code $C\subset{\mathbb{F}}_{q^{2}}^{n}$, we have that $C^{\perp_{h}}=(C^{\perp})^{q}$, where we consider the component wise power of $q$. In particular, this implies that the Hermitian dual and the Euclidean dual have the same parameters. In this section we show that we may apply similar techniques to the ones used in the previous section to compute the Hermitian hull in some cases. In what follows, as we are working over ${\mathbb{F}}_{q^{2}}$, we change $q$ by $q^{2}$ in the definitions of $S$, $A_{i}^{d}$, for $i=1,2,3$, etc. We show now how the main definitions from the other sections are adapted to the Hermitian case in this section:
  1. (1)
    $S={\mathbb{F}}_{q^{2}}[x_{0},x_{1},x_{2}]$.
  2. (2)
    Projective and affine Reed-Muller codes are defined for $1\leq d\leq m(q^{2}-1)$.
  3. (3)
    $A^{d}$ is defined for $1\leq d\leq 2(q^{2}-1)$. If we have $x_{0}^{a_{0}}x_{1}^{a_{1}}x_{2}^{a_{2}}\in A_{1}^{d}$, then $10\leq a_{1},a_{2}\leq q^{2}-$ 1, and if $x_{0}^{a_{0}}x_{1}^{a_{1}}x_{2}^{a_{2}}\in A_{2}^{d}$, then $10\leq a_{2}\leq q^{2}-$ 1. For Lemma 2.9, we have $10\leq a_{i}\leq q^{2}-$ 1, for $1\leq i\leq 2$.
  4. (4)
    Now $\overline{z}$ is the integer $11\leq\overline{z}\leq q^{2}-$ 1 such that $\overline{z}\equiv z\bmod q^{2}-1$ when $z>0$, and $\overline{z}=0$ otherwise.
@@
The codes’ inclusion $\operatorname{RM}_{d}(q^{2},m)\subset\operatorname{RM}_{d}^{\perp_{h}}(q^{2},m)$ holds if, and only if, $10\leq d\leq m(q-1)-$ 1.
@@
The image by the evaluation map over $\mathbb{A}^{2}$ of $U_{d_{1},d_{2}}$ is a basis for $\operatorname{RM}_{d_{1}}(q^{2},2)\cap\operatorname{RM}_{d_{2}}^{\perp_{h}}(q^{2},2)$.
@@
The monomials $x_{1}^{a_{1}}x_{2}^{a_{2}}$ with $10\leq a_{1},a_{2}\leq q^{2}-$ 1 have linearly independent evaluations over $\mathbb{A}^{2}$, and their evaluations generate ${\mathbb{F}}_{q^{2}}^{q^{4}}$ (the full code). The evaluation of a monomial $x_{1}^{a_{1}}x_{2}^{a_{2}}$ with $10\leq a_{1},a_{2}\leq q^{2}-$ 1 and $a_{1}+a_{2}\leq d_{1}$ is in $\operatorname{RM}_{d_{2}}^{\perp_{h}}(q^{2},2)$ if and only if $x_{1}^{a_{1}}x_{2}^{a_{2}}\equiv(x_{1}^{b_{1}}x_{2}^{b_{2}})^{q}\bmod I(\mathbb{A}^{2})$ for some $10\leq b_{1},b_{2}\leq q^{2}-$ 1 such that $b_{1}+b_{2}\leq 2(q^{2}-1)-d_{2}-1$, where $I(\mathbb{A}^{2})=\langle x_{1}^{q^{2}}-x_{1},x_{2}^{q^{2}}-x_{2}\rangle$ (we have used the duality from Theorem 2.5 and the fact that $\operatorname{RM}_{d_{2}}^{\perp_{h}}(q^{2},2)=(\operatorname{RM}_{d_{2}}^{\perp}(q^{2},2))^{q}$). If $a_{i}\neq 0$ for some $i=1,2$, then $x_{1}^{a_{1}}x_{2}^{a_{2}}\equiv(x_{1}^{b_{1}}x_{2}^{b_{2}})^{q}\bmod I(\mathbb{A}^{2})$ implies $a_{i}\equiv qb_{i}\bmod q^{2}-1$ with $b_{i}\neq 0$, which is equivalent to having $b_{i}=\overline{qa_{i}}$ (recall Remark 2.11). If $a_{i}=0$ for some $i=1,2$, then $x_{1}^{a_{1}}x_{2}^{a_{2}}\equiv(x_{1}^{b_{1}}x_{2}^{b_{2}})^{q}\bmod I(\mathbb{A}^{2})$ implies $b_{i}=0=\overline{qa_{i}}$ in this case as well. Therefore, in both cases $b_{i}=\overline{qa_{i}}$, which finishes the proof. ∎
@@
Using that $C^{\perp_{h}}=(C^{\perp})^{q}$, if we consider $(A_{i}^{d})^{q}:=\{(x^{\alpha})^{q}\mid x^{\alpha}\in A_{i}^{d}\}$ and $d^{\perp}=2(q^{2}-1)-d$, we have that the image by the evaluation map of $(A^{d^{\perp}})^{q}:=\bigcup_{i=1}^{3}(A_{i}^{d^{\perp}})^{q}$ is a basis for $\operatorname{PRM}_{d}^{\perp_{h}}(q^{2},2)=(\operatorname{PRM}_{d^{\perp}}(q^{2},2))^{q}$ (if $d\not\equiv 0\bmod q^{2}-1$). Following the notation from the previous section, we will denote by $S_{d^{\perp}}^{q}/I(P^{2})$ the vector space generated in $S/I(P^{2})$ by $\bigcup_{i=1}^{3}(A_{i}^{d^{\perp}})^{q}$.
@@
$\operatorname{PRM}_{d}(q^{2},2)\cap\operatorname{PRM}_{d^{\perp}}^{q}(q^{2},2)=(\operatorname{PRM}_{d}^{q}(q^{2},2)\cap\operatorname{PRM}_{d^{\perp}}(q^{2},2))^{q},$
and then at the right-hand side of the previous equality we have the Hermitian hull of a projective Reed-Muller code of degree $d^{\perp}<q^{2}-1$, to the power of $q$. Moreover, because of Theorem 2.2 we are going to avoid the case with $d=q^{2}-1$ when giving results for the Hermitian hull (for results about $S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$ we will still consider $d=q^{2}-1$). This is because in that case $\operatorname{PRM}_{d}^{\perp}(q^{2},2)=\operatorname{PRM}_{d^{\perp}}(q^{2},2)+\langle(1,\dots,1)\rangle\neq\operatorname{PRM}_{d^{\perp}}(q^{2},2)$, and $\operatorname{PRM}_{d}^{\perp_{h}}(q^{2},2)=(\operatorname{PRM}_{d^{\perp}}(q^{2},2)+\langle(1,\dots,1)\rangle)^{q}\neq\operatorname{PRM}_{d^{\perp}}^{q}(q^{2},2)\cong S_{d^{\perp}}^{q}/I(P^{2})$. Hence, if $d=q^{2}-1$ we do not have the isomorphism between $\operatorname{PRM}_{d}^{\perp_{h}}(q^{2},2)$ and $S_{d^{\perp}}^{q}/I(P^{2})$, and also this case is the least interesting for quantum codes because we do not have a bound for the minimum distance of the dual code.
@@
In the following results we argue in a similar way to Section 3.1 to show which polynomials can be in $S_{d}/I(P^{2})\cap S_{d^{\perp}}^{q}/I(P^{2})$ depending on whether the monomials in the support of these polynomials are contained in $A_{2}^{d}$ or if they have $x_{2}^{d}$ in their support (we recall that $A_{3}^{d}=\{x_{2}^{d}\}$). We restrict to the case $1\leq d\leq 2(q-1)$ in some results because in that case we have $U=A_{1}^{d}$ by Lemma 3.17, which is similar to what happens in the Euclidean case. For the following results, recall that $\overline{z}$ is a representative of the class of $z$ modulo $q^{2}-1$.